OpenStudy (anonymous):

The ratio of two cubes is 216 to 27. What is the ratio of their surface areas?

OpenStudy (radar):

\[216^{2/3}\to 27^{2/3}=4 \to 1\] The cube root brings the measurement to the linear dimensions, the squaring takes it back up to the area unit.

OpenStudy (radar):

\[\sqrt[3]{216}=6\]\[6^{2}=36\]\[\sqrt[3]{27}=\] \[3^{2}=9\] 36:9 is the same as 4:1

OpenStudy (radar):

Put a 3 after the equal sign after cube root 27.

OpenStudy (anonymous):

Thanks!

OpenStudy (radar):

You are welcome.